The value of ∮r/r^3.da on a spherical surfacer r = R is 

(a) 4ΠR2  

(b) Zero

(c) 4Π

(d) infinity


1
Expert's answer
2020-09-07T08:55:40-0400

One needs to evaluate surface integral "\oint_A \frac{\bold r}{r^3} d\bold S = \oint_A \frac{\bold r}{r^3} \bold n d S", where A is the surface of the sphere of radius "R".

Let us use spherical coordinates "\theta, \varphi". The normal vector to the surface of the sphere at point "\bold r" is simply "\frac{\bold r}{r}" , and the surface element is "d S = R^2 \sin \theta d\theta d\varphi".

"\oint_A \frac{\bold r}{r^3} \bold n d S= \int_0^{2\pi}d\varphi\int_0^\pi \frac{r^2}{r^4}|_{r=R} \cdot R^2 \sin \theta d\theta = 2\pi \cdot (-\cos \theta)|_0^{\pi} = 4 \pi".

Hence, the answer is c)